40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15983     Entry time: Thu Apr 1 00:50:06 2021
Author: gautam 
Type: Update 
Category: SUS 
Subject: PRMdiag 

I spent some time investigating the PRM this evening, trying out some of the stuff we discussed in the meeting.

  1. I used one of the SUS lockin oscillator to drive the butterfly mode (UL=+1, UR=-1, LL=-1, LR=+1) of the optic, @4.45 Hz, Amplitude=450cts (Oplev loops were engaged during the measurement).
  2. Used the sensing matrix infrastructure to demodulate the Oplev PIT/YAW error signals, using the other lockin channel (so that oscillator is just for demodulation, it doesn't actuate on the suspension).
  3. The digital demod phase was adjusted to put all of the demodulated signal in one ("I") quadrature.
  4. The balancing of the coils was perturbed by adding small amounts of the naive PIT (YAW) DoF to the pringle-mode actuation, while simultaneously looking to minimize the demodulated signal in YAW (PIT). 

Basically, my finding tonight was that I could not improve (make the pringle mode actuation witnessed by the Oplev QPD smaller) by +/- perturbing the butterfly actuation with of 0.05%, 0.5% and 1% of PIT (I didn't try YAW, or other values of PIT, as none of these seemed to do any good). It seems highly unlikely that the existing coil gains (these come after the output matrix) and the actual coil/magnet pairs are so perfectly tuned, so there must be something wrong with my method. I'll try more combos tomorrow. Separately, I verified that the naive PIT (YAW) moves the optic mainly, i.e. to the eye), in PIT(YAW) as judged by the REFL spot on the camera and the readback of the Oplev QPD.

For this work, I made a few changes to filter banks:

  1. Added 2Hz wide BPs centered around 4.45 Hz to the "SIGNAL" FM of the BS and PRM suspension lockins.
  2. Added 100mHz LPFs to the I and Q demodulated outputs.
  3. Made a copy of Kiwamu's perturbcoilbalance_fourosem.py (in scripts/SUS) to add little bit of PIT/YAW to the pringle actuation.

I noticed that the filters/switch states/gains for LOCKIN1 and LOCKIN2 are not consistent within either PRM or BS suspension, or across suspensions. Several filter INs/OUTs were also disabled - something for the SUSdiag team to note, whenever this is scripted, the script should check that the signal is indeed making it end-to-end.

ELOG V3.1.3-